# Understanding and Mitigating Language Confusion in LLMs

### v1

@@ A.3 Cross-lingual prompt generation
A.4 Metric Variability
@@
A.5 Dataset Language Confusion Metrics
Avg Okapi Aya Dolly Native (Ours)
Llama 2 70B-I 54.0 57.2 49.6 55.2 49.0
Llama 3 70B-I 56.2 56.0 50.8 61.8 47.0
Mixtral 8x7B 74.9 70.7 78.3 75.8 78.2
Mistral Large 75.6 68.4 82.0 76.4 77.8
Command R 98.7 97.8 99.4 99.0 98.8
Command R+ 99.3 99.0 99.0 99.8 99.2
GPT-3.5 Turbo 99.1 99.0 98.6 99.8 99.5
GPT-4 Turbo 99.4 98.9 99.4 99.8 99.2
Table A7: LPR by dataset on monolingual generation.
Monolingual Cross-lingual
Avg Okapi Aya Dolly Native (Ours) Avg Okapi ShareGPT Complex (Ours)
Llama 2 70B-I 97.4 100.0 94.4 97.7 100.0 84.2 90.0 78.9 83.7
Llama 3 70B-I 95.1 100.0 93.3 91.9 90.0 94.4 98.9 87.9 96.4
Mixtral 8x7B 78.0 77.9 74.8 81.3 64.9 68.2 75.8 68.0 60.8
Mistral Large 98.9 100.0 98.9 97.9 99.0 93.8 98.7 95.4 87.3
Command R 96.5 97.4 93.8 98.3 95.5 94.0 95.3 93.7 92.9
Command R+ 99.7 100.0 100.0 99.0 99.5 95.1 97.4 97.6 90.3
GPT-3.5 Turbo 100.0 100.0 100.0 100.0 99.5 98.7 100.0 99.8 96.4
GPT-4 Turbo 99.7 100.0 99.0 100.0 99.5 96.6 99.8 99.5 90.6
Monolingual Cross-lingual
Avg Okapi Aya Dolly Native (Ours) Avg Okapi ShareGPT Complex (Ours)
Llama 2 70B-I 69.4 72.8 65.0 70.5 65.8 52.0 58.6 58.1 39.3
Llama 3 70B-I 70.5 71.8 65.8 73.9 61.8 44.7 51.7 54.5 27.9
Mixtral 8x7B 76.4 74.1 76.5 78.5 71.0 68.3 76.6 73.4 54.8
Mistral Large 85.6 81.2 89.7 85.8 87.1 71.7 80.5 70.8 64.0
Command R 97.6 97.6 96.5 98.7 97.1 76.9 84.3 91.6 54.9
Command R+ 99.5 99.5 99.5 99.4 99.4 93.0 96.7 98.2 84.2
GPT-3.5 Turbo 99.6 99.5 99.3 99.9 99.5 93.8 98.8 98.3 84.4
GPT-4 Turbo 99.5 99.4 99.2 99.9 99.4 93.3 98.2 97.9 83.7
A.6 Extended Impacts
Short Medium Long
[21, 65] [68, 224] [227, 2971]
Command R 42.7 28.1 38.0
   + one-shot 59.5 49.1 67.7
Command R+ 82.1 74.8 79.8
GPT-4 89.3 72.9 71.8
Table A10: Line-level pass rate (LPR) on our “Complex prompts” dataset for crosslingual generation, depending on length of the prompts. We sort the prompts by length and split them into 3 length buckets of the same size (each containing one third of the prompts). [a, b]: min and max length in words of each bucket’s prompts.
Start Integrated End
Command R 86.7 69.0 85.1
   + one-shot 88.7 80.6 90.6
Command R+ 94.4 90.3 95.2
GPT-4 93.0 91.7 95.0
Line-level Pass Rate (LPR)
avg ar hi ja ko ru zh id
FP16 99.3 99.0 100.0 99.0 100.0 100.0 97.5 97.0
W8 99.3 99.7 100.0 98.0 100.0 100.0 99.0 97.0
W8A8 99.5 100.0 100.0 100.0 100.0 100.0 98.5 98.0
W4-g 98.8 99.3 99.0 98.0 100.0 100.0 98.5 91.0
Word-level Pass Rate (WPR)
FP16 98.2 99.3 100.0 100.0 99.0 95.0 96.0 -
W8 98.3 99.3 99.0 99.0 100.0 96.0 96.5 -
W8A8 98.7 99.3 99.0 100.0 100.0 97.0 97.0 -
W4-g 98.1 99.3 100.0 98.0 99.0 93.9 98.5 -
Monolingual Datasets
avg Okapi Aya Dolly Native (Ours)
Llama 2 70B-I 54.0 57.2 49.6 55.2 49.0
Llama 3 70B-I 56.2 56.0 50.8 61.8 47.0
Mixtral 8x7B 74.9 70.7 78.3 75.8 78.2
Mistral Large 75.6 68.4 82.0 76.4 77.8
Command R 98.7 97.8 99.4 99.0 98.8
Command R+ 99.3 99.0 99.0 99.8 99.2
GPT-3.5 Turbo 99.1 99.0 98.6 99.8 99.5
GPT-4 Turbo 99.4 98.9 99.4 99.8 99.2
avg ar hi ja ko ru zh
T=0.0 93.3 98.9 94.6 86.4 96.7 93.0 90.0
T=0.3 94.0 94.3 98.6 88.5 97.2 94.0 91.1
T=0.5 94.2 98.9 96.1 87.1 95.2 95.5 92.3
T=0.7 94.6 97.9 99.1 90.8 95.7 94.6 89.6
T=1.0 80.1 93.5 88.6 70.5 85.5 71.0 71.2
p=0.1 93.6 97.4 96.6 89.0 97.5 93.3 87.6
p=0.3 94.0 98.0 97.6 85.7 98.4 92.0 92.3
p=0.5 94.0 95.0 97.3 89.0 98.8 93.8 90.1
p=0.75 94.0 94.3 98.6 88.5 97.2 94.0 91.1
Table A14: Effect of varying temperature ($T$) or nucleus size ($p$) on cross-lingual word-level language confusion (WPR) of Command R. Default values are $p=0.75$ and $T=0.3$. Best score.
A.7 Quantization
It is common to train at half-precision floating-point (FP16), where weights and activations of a network use 16 bits (2 bytes) to represent a floating-point value. It is common to quantize weights to INT8 (8-bit integers), often referred to as W8. More extreme quantization of weights to INT4 (4-bit) is called W4. Quantizing both weights and activations to INT8 is commonly called W8A8.
In Section 4.4, we compare FP16 with W8, W8A8, and W4242424W4 with group-wise scaling using GPTQ (Frantar et al., 2022). variants of Command R+ on monolingual generation.
A.8 Base vs Instruction Tuning comparison
Avg ar de en es fr hi id it ja ko pt ru tr vi zh
A.9 Nucleus Sampling with Temperature
Previous work shows that greedy decoding empirically leads to repetition and generally low-quality generations. One solution is adding stochasticity to decoding, such as by sampling the next token from the top-K next most likely tokens or from the top-P of the probability distribution over next tokens. The latter is called “nucleus sampling” Holtzman et al. (2019), and can be combined with top-K and temperature sampling.

### v2

@@ A.3 Cross-lingual prompt generation
A.4 Dataset Language Confusion Metrics
Avg Okapi Aya Dolly Native (Ours)
Llama 2 70B-I 54.0 57.2 49.6 55.2 49.0
Llama 3 70B-I 56.2 56.0 50.8 61.8 47.0
Llama 3.1 70B-I 99.7 100.0 100.0 99.3 99.5
Mixtral 8x7B 74.9 70.7 78.3 75.8 78.2
Mistral Large 75.6 68.4 82.0 76.4 77.8
Command R 98.7 97.8 99.4 99.0 98.8
Command R+ 99.3 99.0 99.0 99.8 99.2
Command R Refresh 99.2 98.4 99.0 100.0 99.5
Command R+ Refresh 99.2 99.2 98.6 99.4 99.8
GPT-3.5 Turbo 99.1 99.0 98.6 99.8 99.5
GPT-4 Turbo 99.4 98.9 99.4 99.8 99.2
GPT-4o 99.1 98.8 98.8 99.2 99.5
Monolingual Cross-lingual
Avg Okapi Aya Dolly Native (Ours) Avg Okapi ShareGPT Complex (Ours)
Llama 2 70B-I 97.4 100.0 94.4 97.7 100.0 84.2 90.0 78.9 83.7
Llama 3 70B-I 95.1 100.0 93.3 91.9 90.0 94.4 98.9 87.9 96.4
Llama 3.1 70B-I 99.7 100.0 100.0 99.3 99.5 95.0 97.7 93.3 94.0
Mixtral 8x7B 78.0 77.9 74.8 81.3 64.9 68.2 75.8 68.0 60.8
Mistral Large 98.9 100.0 98.9 97.9 99.0 93.8 98.7 95.4 87.3
Command R 96.5 97.4 93.8 98.3 95.5 94.0 95.3 93.7 92.9
Command R+ 99.7 100.0 100.0 99.0 99.5 95.1 97.4 97.6 90.3
Command R Refresh 99.5 100.0 99.0 99.3 99.5 98.8 99.6 99.6 97.1
Command R+ Refresh 99.6 100.0 99.5 99.0 100.0 98.3 99.3 99.1 96.6
GPT-3.5 Turbo 100.0 100.0 100.0 100.0 99.5 98.7 100.0 99.8 96.4
GPT-4 Turbo 99.7 100.0 99.0 100.0 99.5 96.6 99.8 99.5 90.6
GPT-4o 99.8 100.0 100.0 99.5 99.7 99.0 100.0 99.6 97.4
Monolingual Cross-lingual
Avg Okapi Aya Dolly Native (Ours) Avg Okapi ShareGPT Complex (Ours)
Llama 2 70B-I 69.4 72.8 65.0 70.5 65.8 52.0 58.6 58.1 39.3
Llama 3 70B-I 70.5 71.8 65.8 73.9 61.8 44.7 51.7 54.5 27.9
Llama 3.1 70B-I 99.4 99.5 99.3 99.4 99.3 87.3 91.1 92.1 78.7
Mixtral 8x7B 76.4 74.1 76.5 78.5 71.0 68.3 76.6 73.4 54.8
Mistral Large 85.6 81.2 89.7 85.8 87.1 71.7 80.5 70.8 64.0
Command R 97.6 97.6 96.5 98.7 97.1 76.9 84.3 91.6 54.9
Command R+ 99.5 99.5 99.5 99.4 99.4 93.0 96.7 98.2 84.2
Command R Refresh 99.4 99.2 99.0 99.7 99.5 95.7 98.9 99.5 88.7
Command R+ Refresh 99.4 99.6 99.0 99.2 99.9 96.8 98.6 98.9 93.0
GPT-3.5 Turbo 99.6 99.5 99.3 99.9 99.5 93.8 98.8 98.3 84.4
GPT-4 Turbo 99.5 99.4 99.2 99.9 99.4 93.3 98.2 97.9 83.7
GPT-4o 99.4 99.4 99.4 99.3 99.6 95.5 98.7 98.6 89.1
A.5 Extended Impacts
Short Medium Long
[21, 65] [68, 224] [227, 2971]
Command R 42.7 28.1 38.0
   + one-shot 59.5 49.1 67.7
Command R+ 82.1 74.8 79.8
GPT-4 89.3 72.9 71.8
Table A10: Line-level pass rate (LPR) on our “Complex prompts” dataset for cross-lingual generation, depending on length of the prompts. We sort the prompts by length and split them into 3 length buckets of the same size (each containing one third of the prompts). [a, b]: min and max length in words of each bucket’s prompts.
Start Integrated End
Command R 86.7 69.0 85.1
   + one-shot 88.7 80.6 90.6
Command R+ 94.4 90.3 95.2
GPT-4 93.0 91.7 95.0
A.6 Quantization
It is common to train at half-precision floating-point (FP16), where weights and activations of a network use 16 bits (2 bytes) to represent a floating-point value. It is common to quantize weights to INT8 (8-bit integers), often referred to as W8. More extreme quantization of weights to INT4 (4-bit) is called W4. Quantizing both weights and activations to INT8 is commonly called W8A8. In Section 4.4, we compare FP16 with W8, W8A8, and W4222222W4 with group-wise scaling using GPTQ (Frantar et al., 2022). variants of Command R+ on monolingual generation.
Line-level Pass Rate (LPR)
avg ar hi ja ko ru zh id
FP16 99.3 99.0 100.0 99.0 100.0 100.0 97.5 97.0
W8 99.3 99.7 100.0 98.0 100.0 100.0 99.0 97.0
W8A8 99.5 100.0 100.0 100.0 100.0 100.0 98.5 98.0
W4-g 98.8 99.3 99.0 98.0 100.0 100.0 98.5 91.0
Word-level Pass Rate (WPR)
FP16 98.2 99.3 100.0 100.0 99.0 95.0 96.0 -
W8 98.3 99.3 99.0 99.0 100.0 96.0 96.5 -
W8A8 98.7 99.3 99.0 100.0 100.0 97.0 97.0 -
W4-g 98.1 99.3 100.0 98.0 99.0 93.9 98.5 -
A.7 Extended Beam Search and Nucleus Sampling Results
avg ar hi ja ko ru zh
T=0.0 93.3 98.9 94.6 86.4 96.7 93.0 90.0
T=0.3 94.0 94.3 98.6 88.5 97.2 94.0 91.1
T=0.5 94.2 98.9 96.1 87.1 95.2 95.5 92.3
T=0.7 94.6 97.9 99.1 90.8 95.7 94.6 89.6
T=1.0 80.1 93.5 88.6 70.5 85.5 71.0 71.2
p=0.1 93.6 97.4 96.6 89.0 97.5 93.3 87.6
p=0.3 94.0 98.0 97.6 85.7 98.4 92.0 92.3
p=0.5 94.0 95.0 97.3 89.0 98.8 93.8 90.1
p=0.75 94.0 94.3 98.6 88.5 97.2 94.0 91.1
Table A13: Effect of varying temperature ($T$) or nucleus size ($p$) on cross-lingual word-level language confusion (WPR) of Command R. Default values are $p=0.75$ and $T=0.3$. Best score is in bold.
Monolingual
avg ar hi ja ko ru zh
1 97.8 98.6 100.0 97.0 97.0 98.0 96.4
2 98.6 98.2 100.0 98.0 100.0 98.0 97.4
3 98.6 98.9 100.0 100.0 98.0 98.0 96.9
5 99.0 99.3 100.0 99.0 98.0 100.0 97.4
10 99.0 98.5 100.0 99.0 98.0 100.0 98.4
Cross-lingual
avg ar hi ja ko ru zh
1 94.9 98.9 97.2 89.4 97.5 95.8 90.4
2 95.4 99.2 97.1 91.1 95.7 97.2 92.4
3 97.1 100.0 97.9 95.7 98.0 96.4 94.5
5 96.7 99.7 98.3 95.1 96.4 96.2 94.7
10 96.7 98.3 98.3 96.7 98.4 94.5 93.7
Table A14: Monolingual and Cross-lingual word-level pass rate (WPR) of Command R using beam search with beam sizes 1–10. ($\neg$) IE = (non-) Indo-European language. ($\neg$) Latin = (non-) Latin script.
Monolingual
avg ar de en es fr hi id it ja ko pt ru tr vi zh $\neg$ IE IE $\neg$ Latin Latin
T=0.0 98.8 100.0 98.0 99.0 96.7 99.0 100.0 96.0 99.0 99.0 100.0 98.0 100.0 99.0 100.0 99.0 99.1 98.5 99.7 98.3
T=0.3 98.6 100.0 98.0 99.5 95.7 99.3 100.0 92.0 99.0 100.0 100.0 98.5 100.0 99.0 99.0 98.5 98.6 98.6 99.7 97.8
T=0.5 98.7 99.0 98.0 99.0 96.3 99.7 100.0 94.0 99.0 100.0 100.0 98.5 100.0 99.0 99.0 98.5 98.7 98.6 99.6 98.1
T=0.7 98.9 100.0 98.0 99.0 95.7 99.0 99.0 99.0 100.0 99.0 100.0 99.0 100.0 99.0 99.0 98.5 99.3 98.5 99.4 98.6
T=1.0 96.8 100.0 96.0 99.5 95.0 99.7 98.0 91.9 98.0 99.0 98.0 97.0 97.0 92.9 93.0 97.5 96.2 97.6 98.2 95.9
p=0.1 98.9 100.0 98.0 99.0 96.7 99.0 100.0 96.0 99.0 100.0 100.0 98.5 100.0 99.0 100.0 99.0 99.2 98.6 99.8 98.3
p=0.3 98.9 100.0 98.0 99.0 96.3 99.0 100.0 96.0 99.0 99.0 100.0 98.5 100.0 99.0 100.0 99.0 99.1 98.5 99.7 98.3
p=0.5 98.8 100.0 98.0 99.0 96.3 98.7 100.0 96.0 99.0 99.0 100.0 98.5 100.0 99.0 100.0 98.5 99.1 98.5 99.6 98.3
p=0.75 98.6 100.0 98.0 99.5 95.7 99.3 100.0 92.0 99.0 100.0 100.0 98.5 100.0 99.0 99.0 98.5 98.6 98.6 99.7 97.8
Cross-lingual
avg ar de en es fr hi id it ja ko pt ru tr vi zh $\neg$ IE IE $\neg$ Latin Latin
T=0.0 68.1 61.7 61.9 - 70.7 74.6 65.2 70.4 65.9 68.1 70.8 68.8 68.8 65.6 65.7 74.8 68.2 67.8 68.2 67.9
T=0.3 68.1 61.6 63.2 - 72.5 74.4 65.5 70.8 65.7 65.3 69.2 67.2 69.4 67.7 65.7 75.0 68.1 68.1 67.7 68.4
T=0.5 67.9 60.7 62.5 - 72.1 74.3 64.3 69.6 66.2 67.1 70.2 67.8 67.1 67.0 65.6 76.2 67.9 67.9 67.6 68.1
T=0.7 67.9 60.7 63.2 - 72.1 74.3 65.6 68.8 66.2 68.1 69.7 68.4 68.4 64.4 65.7 75.5 67.7 68.3 68.0 67.9
T=1.0 64.0 63.3 63.0 - 70.6 72.4 63.3 63.9 62.0 65.4 65.2 66.1 62.7 57.1 51.7 69.8 62.4 66.2 64.9 63.3
p=0.1 68.0 61.3 62.5 - 71.1 75.0 65.1 70.1 66.9 67.8 70.2 68.4 68.3 65.9 65.7 74.2 67.9 68.2 67.8 68.2
p=0.3 68.1 61.3 62.5 - 70.7 74.6 65.4 69.7 66.6 67.7 69.8 68.1 68.4 67.2 65.7 75.2 68.1 68.0 68.0 68.1
p=0.5 68.2 61.1 62.9 - 71.4 75.3 65.1 69.2 66.6 68.1 70.1 68.8 69.5 66.8 65.9 74.5 68.2 68.3 68.1 68.4
p=0.75 68.1 61.6 63.2 - 72.5 74.4 65.5 70.8 65.7 65.3 69.2 67.2 69.4 67.7 65.7 75.0 68.1 68.1 67.7 68.4
Table A15: Effect of varying temperature ($T$) or nucleus size ($p$) on monolingual and crosslingual line-level language confusion (LPR) of Command R. Default values are $p=0.75$ and $T=0.3$. Best score. ($\neg$) IE = (non-) Indo-European language. ($\neg$) Latin = (non-) Latin script.
Monolingual
avg ar de en es fr hi id it ja ko pt ru tr vi zh $\neg$ IE IE $\neg$ Latin Latin
1 99.0 100.0 98.0 99.5 96.0 99.3 100.0 97.0 99.0 100.0 100.0 99.5 100.0 97.9 100.0 99.0 99.2 98.8 99.8 98.5
2 99.0 99.6 98.0 99.5 95.6 99.7 100.0 96.0 99.0 100.0 100.0 99.0 100.0 98.9 100.0 99.0 99.2 98.7 99.8 98.4
3 98.7 99.6 98.0 99.0 95.7 99.7 99.0 96.0 99.0 99.0 100.0 99.0 100.0 98.9 100.0 98.0 98.9 98.5 99.3 98.3
5 99.0 99.3 98.0 99.5 96.0 99.7 100.0 96.0 100.0 100.0 100.0 99.5 100.0 98.9 100.0 98.5 99.1 98.9 99.6 98.6
10 98.5 98.9 98.0 99.5 95.7 99.3 100.0 96.0 100.0 99.0 100.0 98.0 100.0 97.8 99.0 96.9 98.4 98.6 99.1 98.1
Cross-lingual
avg ar de en es fr hi id it ja ko pt ru tr vi zh $\neg$ IE IE $\neg$ Latin Latin
1 74.1 68.4 70.9 - 76.9 77.2 74.7 74.9 72.0 73.5 76.4 73.7 73.6 73.8 73.0 77.8 73.9 74.2 74.1 74.0
2 72.2 67.0 70.5 - 76.5 75.7 74.0 72.5 70.8 70.9 73.8 74.3 70.8 69.5 70.8 73.5 71.1 73.6 71.6 72.6
3 71.5 67.2 69.8 - 76.5 76.4 75.0 69.8 71.6 69.2 73.1 71.1 69.3 67.6 70.1 74.8 70.1 73.4 71.4 71.6
5 70.3 65.0 69.6 - 76.2 76.3 75.1 67.0 71.0 68.2 72.4 69.1 67.6 65.5 66.7 73.8 68.3 72.9 70.4 70.2
10 68.4 65.0 68.6 - 75.9 75.7 75.7 66.2 70.2 66.1 68.8 66.7 63.7 60.6 64.7 69.6 65.6 72.1 68.1 68.6
Table A16: Monolingual and Cross-lingual line-level pass rate (LPR) of Command R using beam search with beam sizes 1-10. ($\neg$) IE = (non-) Indo-European language. ($\neg$) Latin = (non-) Latin script.
A.8 Base vs Instruction Tuning comparison
Avg ar de en es fr hi id it ja ko pt ru tr vi zh
Model Crosslingual WPR Crosslingual LPR
avg ar hi ja ko ru zh avg ar de es fr hi id it ja ko pt ru tr vi zh
Command R Base 100.0 100.0 100.0 100.0 100.0 100.0 100.0 1.1 0.0 0.9 1.2 1.6 0.8 1.4 0.3 2.4 0.0 0.9 0.0 0.4 1.6 3.6
+ Q/A template 97.0 95.2 100.0 99.0 97.9 91.7 98.4 20.9 3.3 19.9 27.6 27.5 5.3 25.8 23.2 42.5 13.5 17.5 8.0 19.6 14.9 44.0
+ 1-shot 98.6 98.5 98.8 98.6 98.3 98.4 99.2 90.7 83.8 93.3 95.0 93.2 84.2 86.7 93.0 93.2 92.5 91.1 88.3 89.6 91.9 93.3
+ 5-shot 99.7 100.0 99.2 100.0 99.5 100.0 99.6 95.0 91.0 96.3 96.8 95.9 96.0 92.8 96.5 93.8 95.5 95.4 93.8 96.8 95.6 94.5
+ English SFT 91.7 95.5 98.7 77.9 94.8 94.3 89.0 78.3 70.0 77.7 82.7 76.6 78.4 77.7 74.8 80.8 80.7 79.3 74.4 82.4 83.9 77.1
   + English pref. tuning 87.4 91.2 97.7 70.4 92.9 88.5 83.6 85.7 86.4 87.4 86.2 82.1 87.7 80.2 83.9 87.6 84.3 83.2 86.7 90.1 89.6 84.6
+ Multilingual SFT 90.0 96.0 97.6 76.0 92.1 89.0 89.7 78.2 89.1 62.4 76.8 69.4 82.8 77.3 72.6 77.7 81.1 76.8 79.5 82.3 85.1 82.4
   + Multi. pref. tuning 86.9 94.1 96.6 71.2 87.7 84.9 87.0 89.4 94.1 80.9 91.9 88.7 91.2 85.6 88.9 89.2 93.0 83.7 92.3 91.7 93.0 87.8
Command R 94.0 94.3 98.6 88.5 97.2 94.0 91.1 68.1 61.6 63.2 72.5 74.4 65.5 70.8 65.7 65.3 69.2 67.2 69.4 67.7 65.7 75.0
+ 1-shot 92.3 97.2 98.3 87.6 89.7 91.6 89.5 82.9 80.0 79.5 85.6 82.6 84.7 79.9 82.4 78.0 87.2 81.5 84.4 84.5 84.5 85.2
Table A18: Effect of few-shot prompting and instruction tuning on cross-lingual language confusion, detailed per-language results.
Model Monolingual WPR Monolingual LPR
avg ar hi ja ko ru zh avg ar de en es fr hi id it ja ko pt ru tr vi zh
Command R Base 98.7 97.9 95.1 100.0 100.0 100.0 99.4 86.2 94.9 85.0 99.5 90.5 85.3 81.0 72.2 74.0 93.9 94.2 84.0 94.8 92.0 83.0 68.1
+ Q/A template 99.7 100.0 100.0 100.0 100.0 98.5 100.0 85.3 87.6 73.2 100.0 89.8 86.3 98.9 49.0 91.9 97.9 93.5 96.8 86.7 96.3 49.0 82.4
+ 1-shot 100.0 100.0 100.0 100.0 100.0 100.0 100.0 94.1 93.0 99.0 100.0 99.3 47.5 96.9 93.8 99.0 99.0 96.8 99.0 97.3 95.1 96.6 99.0
+ 5-shot 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.0 99.6 99.0 100.0 100.0 98.3 98.0 96.9 100.0 100.0 99.0 99.5 98.9 98.8 99.0 97.9
+ English SFT 96.2 98.4 100.0 88.1 98.5 100.0 92.3 77.8 54.1 83.8 100.0 80.3 86.9 59.6 89.0 96.0 67.7 66.7 84.4 68.0 64.0 85.0 82.0
   + English pref. tuning 90.9 93.5 97.8 84.4 95.4 89.6 85.0 74.3 54.4 81.0 100.0 82.7 85.6 46.0 80.0 95.0 64.0 65.0 81.4 67.7 51.0 82.0 78.5
+ Multilingual SFT 95.5 97.9 100.0 86.6 97.9 99.0 91.6 98.3 99.7 100.0 100.0 99.7 98.7 100.0 90.0 99.0 98.0 98.0 98.4 100.0 96.8 100.0 97.0
   + Multi. pref. tuning 93.4 97.9 100.0 86.7 93.0 91.8 91.2 98.8 99.6 97.0 99.5 99.3 99.3 100.0 92.0 100.0 99.0 100.0 99.0 100.0 100.0 100.0 98.0
Command R 96.3 99.3 99.0 93.9 97.0 96.0 92.3 98.6 100.0 98.0 99.5 95.7 99.3 100.0 92.0 99.0 100.0 100.0 98.5 100.0 99.0 99.0 98.5
+ 1-shot 92.7 94.4 95.7 93.8 88.9 90.5 92.9 68.3 51.8 53.0 5.5 83.2 98.3 23.7 88.0 94.0 65.0 95.7 66.0 95.5 90.3 97.8 17.0
Table A19: Effect of few-shot prompting and instruction tuning on monolingual language confusion, detailed per-language results.
A.9 Metric Variability
@@
A.10 Nucleus Sampling with Temperature
Previous work shows that greedy search empirically leads to repetition and generally low-quality generations. One solution is adding stochasticity to decoding, such as by sampling the next token from the top-K next most likely tokens or from the top-P of the probability distribution over next tokens. The latter is called “nucleus sampling” Holtzman et al. (2019), and can be combined with top-K and temperature sampling.